# Thermal evolution of dark matter in the early universe from a symplectic glueball model

### v1

@@ 3 Results
It is interesting to note that, in the temperature range under consideration, the quantities plotted in figure 4 are still far from their Stefan–Boltzmann-limit values (which are denoted by the horizontal bars on the right-hand side of the plot. This shows that, for temperatures up to a few times $T_{\mbox{\tiny{c}}}$, the deconfined state of matter can hardly be described as an ideal (or nearly ideal) gas of gluon-like particles: instead, the clearly non-vanishing values for the trace of the energy-momentum tensor reveal that interactions among the degrees of freedom present in the deconfined phase are non-negligible. Also, when the temperature is increased, the approach towards the Stefan–Boltzmann limit is expected to be slow: in particular, if one takes the temperature to be the characteristic (hard) momentum scale in the deconfined phase, then eq. (3.5) shows that, at leading order, the coupling depends logarithmically on the temperature, i.e., it decreases very slowly as a function of $T$. These features are common to other non-Abelian gauge theories.
@@
which can be immediately evaluated from our extrapolated results for the entropy density. The results are shown in the main plot of figure 5, where the displayed uncertainty band is dominated by the systematic uncertainties affecting these data, which we tentatively estimate to be of the order of $10\%$. For comparison, in the inset plot of figure 5 we also show the analogous results (but without their uncertainties) obtained from lattice simulations of $\operatorname{\mathrm{SU}}(3)$ Yang–Mills theory [54, 178], and recently discussed in ref. [179]. One notes that the behavior of the speed of sound in the two theories is qualitatively, and even quantitatively, very similar.
Figure 5: Squared speed of sound in the $\operatorname{\mathrm{Sp}}(2)$ theory, plotted against $T/T_{\mbox{\tiny{c}}}$; the displayed error band includes both statistical uncertainties and a (crude, and somewhat conservative) estimate of the systematic uncertainties related to the continuum extrapolation of our results. The horizontal bar on the right-hand side of the plot denotes the value of squared speed of sound in the Stefan–Boltzmann limit, which is $c_{\mbox{\tiny{s}}}^{2}=\frac{1}{3}$. The inset plot (in which, for the sake of clarity, the uncertainties are not displayed) shows the corresponding results that can be obtained from the lattice calculations of $\operatorname{\mathrm{SU}}(3)$ Yang–Mills theory in the deconfined phase [54] and in the confining phase [178], as discussed in ref. [179].
Another feature of the $\operatorname{\mathrm{Sp}}(2)$ equation of state, which is shared with gauge theories based on other non-Abelian groups, is the fact that, at low temperatures in the confining phase, the equilibrium thermodynamic quantities can be accurately modelled in terms of a gas of massive and non-interacting glueballs [180, 181, 182, 183, 184] (see also ref. [185] for a discussion from a modern perspective). For example, the pressure can be written as the sum of the contributions from the different glueball states, labelled by $k$, with masses $m_{k}$:
@@
Using a known asymptotic expansion of Bessel functions [186], at large $\frac{m_{k}}{T}$ eq. (3.8) can be rewritten as
@@
The comparison of our extrapolated simulation results with the prediction from eq. (3.8) is shown in figure 6, where the dash-dotted line denotes the dominating contribution, from the lightest glueball, to the $p/T^{4}$ ratio, while the dashed line includes the contributions from all glueballs below the two-particle threshold. As the plot shows, the continuum-extrapolated results obtained from our lattice calculations (denoted by the solid red curve) are consistent with the prediction from the glueball-gas model at low temperatures, but for $T\gtrsim 0.9T_{\mbox{\tiny{c}}}$ the curve obtained from eq. (3.8) underestimates the lattice results. As in $\operatorname{\mathrm{SU}}(3)$ Yang–Mills theory [187, 54], this behavior can be interpreted in terms of a continuous, exponentially growing, Hagedorn-like, density of states.
@@
For each value of $N_{s}$, we determined the maximum value of the specific heat per volume from the lattice results for $(C_{V}/V)a_{\mbox{\tiny{lat}}}^{6}$ at different values of $\beta$, by interpolating the datum corresponding to the largest value of $C_{V}$, the two data points at its left, and the two at its right, through a quadratic fit in $\beta$, and extracting the maximum value of the fitted parabola. In figure 8 we show our study of the latent heat of the theory (from simulations with $N_{\tau}=5$), based on eq. (2.21): the plot shows the extrapolation of the peak values for the specific heat to the infinite-volume limit according to
@@
The fit of the data shown in figure 8 yields
$\gamma_{0}=1.73(32)\cdot 10^{-4},\qquad\gamma_{1}=8.53(67),\qquad\gamma_{2}=-5.0(3.4)\cdot 10^{3},$ (3.11)
with $\chi^{2}_{\mbox{\tiny{red}}}=0.09$. The smallness of the reduced $\chi^{2}$ and the fact that the fit result for $\gamma_{2}$ is compatible with zero within less than one standard deviation and a half indicates that a quadratic fit in $1/(Va_{\mbox{\tiny{lat}}}^{3})$ is an overfitting. For comparison, a two-parameter fit with $\gamma_{2}$ set to zero, leads to
$\gamma_{0}=2.26(20)\cdot 10^{-4},\qquad\gamma_{1}=7.40(24),$ (3.12)
with $\chi^{2}_{\mbox{\tiny{red}}}=0.67$. Assuming the extrapolation according to this linear fit, and including its discrepancy with respect to the quadratic fit as a systematic uncertainty, the latent heat obtained from our $N_{\tau}=5$ simulations (and expressed in units of the fourth power of the inverse lattice spacing corresponding to the critical temperature for this lattice) is found to be consistent with the one that can be determined from the discontinuity in the energy density of the theory across the deconfinement temperature, according to the procedure described in refs. [59, 188], which in our case leads to $\frac{L_{\mbox{\tiny{h}}}}{T_{\mbox{\tiny{c}}}^{4}}=1.76(9)(11)$ for $N_{\tau}=5$ and to $\frac{L_{\mbox{\tiny{h}}}}{T_{\mbox{\tiny{c}}}^{4}}=1.50(15)(17)$ for $N_{\tau}=6$ (where the first error is statistical, whereas the second accounts for systematic uncertainties, including, in particular, those associated with varying the number of data used in the infinite-volume extrapolation, and with performing such extrapolation including or excluding, beside a correction linear in the inverse volume, also a quadratic one): within their uncertainties (which can be summed in quadrature), both results are consistent with the approximate continuum-extrapolated value
@@
This value can be compared with analogous results in $\operatorname{\mathrm{SU}}(N)$ Yang–Mills theories. Excluding the $\operatorname{\mathrm{SU}}(2)$ theory, which has a second-order deconfinement transition [159, 189, 190], and hence a vanishing latent heat, the $\operatorname{\mathrm{SU}}(3)$ theory has a relatively weak first-order transition, with $\frac{L_{\mbox{\tiny{h}}}}{T_{\mbox{\tiny{c}}}^{4}}\simeq 1.4$ [191], while the values of the latent heat observed in $\operatorname{\mathrm{SU}}(N)$ gauge theories with more than three color charges scale well with the dimension of the gauge group and are described by $\frac{L_{\mbox{\tiny{h}}}}{N^{2}T_{\mbox{\tiny{c}}}^{4}}\simeq 0.344$ [188]. For the $\operatorname{\mathrm{Sp}}(2)$ theory, the dimension of the gauge group is ten, i.e., intermediate between those of $\operatorname{\mathrm{SU}}(3)$ and $\operatorname{\mathrm{SU}}(4)$, and a discontinuous deconfinement transition is observed; hence, if one would argue that this Yang–Mills theory should behave similarly to those based on unitary gauge groups, and would simply attempt crude estimate of the $\frac{L_{\mbox{\tiny{h}}}}{T_{\mbox{\tiny{c}}}^{4}}$ ratio by rescaling those found for the $\operatorname{\mathrm{SU}}(3)$ and $\operatorname{\mathrm{SU}}(4)$ theories according to the number of gluon degrees of freedom, one would get a result in the ballpark between approximately $1.5$ and $3$. The estimate (3.13) obtained from our simulation results lies within this interval.
@@
Figure 8: Extrapolation of our lattice results at $N_{\tau}=5$ for the maximum value of the specific heat divided by the volume and expressed in the appropriate units of the lattice spacing (blue circles) to their infinite-volume limit. The dashed green curve is the result of a parabolic fit according to eq. (3.10), while the solid red straight line is obtained by a linear fit of the data (i.e., setting $\gamma_{2}$ to zero). The corresponding extrapolated values, respectively denoted by the green triangle and by the red square, are reported in eq. (3.11) and in eq. (3.12), and are compatible with each other.
@@
One of the reasons why it was interesting for us to study the finite-temperature confining transition in a symplectic Yang–Mills theory consisted in investigating, whether the center of the gauge group, which is widely believed to be directly responsible for the emergence of confinement at large distances in the ground state of non-Abelian gauge theories, via the formation of center vortices [140, 141, 142, 143], also affects the nature of the confinement/deconfinement transition at finite temperature. In particular, general arguments based on the renormalization group indicate that the thermal deconfinement transition for $\operatorname{\mathrm{Sp}}(N)$ gauge theories could be of the second order, and in the same universality as the three-dimensional Ising model [57]. Our results, which are consistent with previous lattice studies of this issue [34], show that this is not the case: the transition is of first order, with a finite latent heat that, when expressed in units of the fourth power of the deconfinement temperature, is of the same order of magnitude as, and slightly larger than, in the $\operatorname{\mathrm{SU}}(3)$ Yang–Mills theory. This confirms the conjecture, that the continuous or discontinuous nature of the confinement/deconfinement transition is essentially determined by the imbalance in the number of degrees of freedom in the high- and in the low-temperature phases, i.e., it depends on the size of the gauge group, not directly on the center symmetry [34]. We also note that, in fact, the discontinuous nature of the deconfinement transition for the $\operatorname{\mathrm{Sp}}(2)$ theory is also supported by analytical arguments, including those based on calculations starting from a supersymmetric version of the theory on a manifold with a compactified spatial direction [192], and those derived within the framework of the function renormalization group [193].
@@
4 Interpretation as a model for dark matter
Before discussing some of the implications predicted from our model for the thermal evolution of dark matter in the early universe, it is important to point out that, if the dark sector is not directly coupled with the SM sector, then its temperature can be different from the temperature of the latter.
If the dark sector is subject to reheating in the early stages of the universe, one expects that it will first be in the deconfined phase at high temperatures, then it will undergo a cosmological phase transition to the confining phase. Thus, the non-Abelian gauge theory describing the dark sector will undergo an evolution qualitatively similar to the one of quantum chromodynamics, from a “hot”, deconfined phase, to a “cold”, confining phase.
At sufficiently high temperatures realized in the early universe, the dark sector that we are studying in the present work consists of a plasma of $\operatorname{\mathrm{Sp}}(2)$ “dark gluons”, which interact with each other through screened long-range forces. Like in a generic non-Abelian gauge theory (see, for example, refs. [196, 197] and references therein), at sufficiently high temperatures a hierarchy of scales

### v2

@@ 3 Results
It is interesting to note that, in the temperature range under consideration, the quantities plotted in figure 5 are still far from their Stefan–Boltzmann-limit values (which are denoted by the horizontal bars on the right-hand side of the plot): this can be clearly seen in the inset plot, displaying $s/T^{3}$, $\epsilon/T^{4}$, and $p/T^{4}$ divided by their Stefan–Boltzmann limit values. This shows that, for temperatures up to a few times $T_{\mbox{\tiny{c}}}$, the deconfined state of matter can hardly be described as an ideal (or nearly ideal) gas of gluon-like particles: instead, the clearly non-vanishing values for the trace of the energy-momentum tensor reveal that interactions among the degrees of freedom present in the deconfined phase are non-negligible. Also, when the temperature is increased, the approach towards the Stefan–Boltzmann limit is expected to be slow: in particular, if one takes the temperature to be the characteristic (hard) momentum scale in the deconfined phase, then eq. (3.5) shows that, at leading order, the coupling depends logarithmically on the temperature, i.e., it decreases very slowly as a function of $T$. These features are common to other non-Abelian gauge theories.
@@
which can be immediately evaluated from our extrapolated results for the entropy density. The results are shown in the main plot of figure 6, where the displayed uncertainty band is dominated by the systematic uncertainties affecting these data, which we tentatively estimate to be of the order of $10\%$. For comparison, in the inset plot of figure 6 we also show the analogous results (but without their uncertainties) obtained from lattice simulations of $\operatorname{\mathrm{SU}}(3)$ Yang–Mills theory [64, 206], and discussed in ref. [207]. One notes that the behavior of the speed of sound in the two theories is qualitatively, and even quantitatively, very similar.
Figure 6: Squared speed of sound in the $\operatorname{\mathrm{Sp}}(2)$ theory, plotted against $T/T_{\mbox{\tiny{c}}}$; the displayed error band includes both statistical uncertainties and a (crude, and somewhat conservative) estimate of the systematic uncertainties related to the continuum extrapolation of our results. The horizontal bar on the right-hand side of the plot denotes the value of squared speed of sound in the Stefan–Boltzmann limit, which is $c_{\mbox{\tiny{s,\,SB}}}^{2}=1/3$, and, in order to help visualize the approach to that limit, the right-hand-side vertical axis shows the values of the $c_{\mbox{\tiny{s}}}^{2}/c_{\mbox{\tiny{s,\,SB}}}^{2}$ ratio. The inset plot (in which, for the sake of clarity, the uncertainties are not displayed) shows the results for $c_{\mbox{\tiny{s}}}^{2}$ that can be obtained from the lattice calculations of $\operatorname{\mathrm{SU}}(3)$ Yang–Mills theory in the deconfined phase [64] and in the confining phase [206], as discussed in ref. [207].
Another feature of the $\operatorname{\mathrm{Sp}}(2)$ equation of state, which is shared with gauge theories based on other non-Abelian groups, is the fact that, at low temperatures in the confining phase, the equilibrium thermodynamic quantities can be accurately modelled in terms of a gas of massive and non-interacting glueballs [208, 209, 210, 211, 212] (see also ref. [213] for a discussion from a modern perspective). For example, the pressure can be written as the sum of the contributions from the different single-particle glueball states, labelled by $k$, with masses $m_{k}$:
@@
Using a known asymptotic expansion of Bessel functions [214], at large $m_{k}/T$ eq. (3.8) can be rewritten as
@@
The comparison of our extrapolated simulation results with the prediction from eq. (3.8) is shown in figure 7, where the dash-dotted line denotes the dominating contribution, from the lightest glueball, to the $p/T^{4}$ ratio, while the dashed line includes the contributions from all glueballs below the two-particle threshold. As the plot shows, the continuum-extrapolated results obtained from our lattice calculations (denoted by the solid red curve) are consistent with the prediction from the glueball-gas model at low temperatures, but for $T\gtrsim 0.9T_{\mbox{\tiny{c}}}$ the curve obtained from eq. (3.8) underestimates the lattice results. As in $\operatorname{\mathrm{SU}}(3)$ Yang–Mills theory [215, 64], this behavior can be interpreted in terms of a continuous, exponentially growing, Hagedorn-like, density of states. Note that, while in principle the inclusion of high-precision results obtained from finer lattices ($N_{\tau}\geq 8$) would be useful to corroborate the continuum extrapolation of the equation of state at low temperatures, in practice the computational costs to reach uncertainties comparable with those of the data evaluated from the other ensembles considered in this work, up to $N_{\tau}=7$, would quickly become prohibitive. The problem is especially severe at $T\leq T_{\mbox{\tiny{c}}}$, since in the confining phase the thermodynamic quantities take values that are significantly smaller than in the deconfined phase, and that, furthermore, are exponentially suppressed when the temperature is reduced. For this reason, in figure 7 we did not include, for example, the results of our simulations on lattices with $N_{\tau}=8$ and $N_{\tau}=9$ at $T=0.9T_{\mbox{\tiny{c}}}$ and at $T=0.95T_{\mbox{\tiny{c}}}$ that are shown in figure 3: they are compatible with zero within their uncertainties, they would not be sufficient to obtain any meaningful interpolating curve for the pressure at either of those $N_{\tau}$ values (for this purpose, the numerical integration would require a much larger number of data points, at sufficiently close temperatures), and would not add any significant information to the curve displayed in figure 7.
@@
For each value of $N_{s}$, we determined the maximum of the specific heat per volume from the lattice results for $(C_{V}/V)a_{\mbox{\tiny{lat}}}^{6}$ at different values of $\beta$, by interpolating the datum corresponding to the largest value of $C_{V}$, the two data points at its left, and the two at its right, through a quadratic fit in $\beta$, and extracting the maximum value of the fitted parabola. In figure 9 we show our study of the latent heat of the theory, based on eq. (2.21): the plot shows the extrapolation of the peak values for the specific heat to the infinite-volume limit according to
@@
For the data extracted from our simulations with $N_{\tau}=5$, the fit of the data shown in figure 9 yields
$\gamma_{0}=3.8(7)\cdot 10^{-5},\qquad\gamma_{1}=1.87(15),\qquad\gamma_{2}=-1.1(7)\cdot 10^{3},$ (3.11)
with $\chi^{2}_{\mbox{\tiny{red}}}=0.09$. The smallness of the reduced $\chi^{2}$ and the fact that the fit result for $\gamma_{2}$ is compatible with zero within approximately one standard deviation and a half indicates that a quadratic fit in $1/(Va_{\mbox{\tiny{lat}}}^{3})$ is an overfitting. For comparison, a two-parameter fit with $\gamma_{2}$ set to zero, leads to
$\gamma_{0}=5.0(4)\cdot 10^{-5},\qquad\gamma_{1}=1.62(5),$ (3.12)
with $\chi^{2}_{\mbox{\tiny{red}}}=0.67$. Assuming the extrapolation according to this linear fit, and including its discrepancy with respect to the quadratic fit as a systematic uncertainty, the latent heat obtained from our $N_{\tau}=5$ simulations (and expressed in units of the fourth power of the inverse lattice spacing corresponding to the critical temperature for this lattice) is found to be consistent with the one that can be determined from the discontinuity in the energy density of the theory across the deconfinement temperature, according to the procedure described in refs. [69, 216], which in our case leads to $L_{\mbox{\tiny{h}}}/T_{\mbox{\tiny{c}}}^{4}=1.76(9)(11)$ for $N_{\tau}=5$, to $L_{\mbox{\tiny{h}}}/T_{\mbox{\tiny{c}}}^{4}=1.5(1)(6)$ for $N_{\tau}=6$ (where the first error is statistical, whereas the second accounts for systematic uncertainties, including, in particular, those associated with varying the number of data used in the infinite-volume extrapolation, and with performing such extrapolation including or excluding, beside a correction linear in the inverse volume, also a quadratic one). Within their uncertainties (which can be summed in quadrature), these results are consistent with the approximate continuum-extrapolated value
@@
As a consistency check, we also carried out a set of computations on finer lattices, with $N_{\tau}=7$. In this case we run simulations for three volumes only (corresponding to $N_{s}=24$, $N_{s}=28$, and $N_{s}=30$) and, as a consequence, performed the thermodynamic extrapolation of the $a_{\mbox{\tiny{lat}}}^{6}C^{\mbox{\tiny{max}}}_{V}/V$ ratio considering only the two-parameter fit with $\gamma_{2}$ set to zero, obtaining $\gamma_{0}=6.5(3.7)\cdot 10^{-6}$ and $\gamma_{1}=0.26(7)$. The corresponding value for the latent heat in units of the fourth power of the deconfinement temperature is $L_{\mbox{\tiny{h}}}/T_{\mbox{\tiny{c}}}^{4}=1.8(5)$: even though the uncertainty is large (due to the fact that, for this $N_{\tau}$ value, the extrapolated $\gamma_{0}$ value has a relative error of the order of $50\%$), the value is compatible with our continuum estimate reported in eq. (3.13). It is important to stress that a reliable estimate of the $C^{\mbox{\tiny{max}}}_{V}/V$ ratio becomes increasingly difficult when $N_{\tau}$ is increased, because this quantity, when evaluated from a lattice calculation, is expressed as a dimensionless number proportional to $a_{\mbox{\tiny{lat}}}^{6}$; given that the statistical uncertainties in a Monte Carlo calculation scale with the inverse of the square root of the number of configurations, this implies that the computational costs to evaluate the $C^{\mbox{\tiny{max}}}_{V}/V$ ratio on the lattice with a fixed relative error are expected to grow at least as $N_{\tau}^{12}$.
@@
The continuum estimate for the $L_{\mbox{\tiny{h}}}/T_{\mbox{\tiny{c}}}^{4}$ in eq. (3.13) can be compared with analogous results in $\operatorname{\mathrm{SU}}(N)$ Yang–Mills theories. Excluding the $\operatorname{\mathrm{SU}}(2)$ theory, which has a second-order deconfinement transition [176, 217, 218], and hence a vanishing latent heat, the $\operatorname{\mathrm{SU}}(3)$ theory has a relatively weak first-order transition, with $L_{\mbox{\tiny{h}}}/T_{\mbox{\tiny{c}}}^{4}\simeq 1.4$ [219], while the values of the latent heat observed in $\operatorname{\mathrm{SU}}(N)$ gauge theories with more than three color charges scale well with the dimension of the gauge group and are described by $L_{\mbox{\tiny{h}}}/(N^{2}T_{\mbox{\tiny{c}}}^{4})=0.360(6)-1.88(17)/N^{2}$ [220] (which is consistent with, and significantly more precise than, the earlier result $L_{\mbox{\tiny{h}}}/(N^{2}T_{\mbox{\tiny{c}}}^{4})\simeq 0.344(72)$ [216]). For the $\operatorname{\mathrm{Sp}}(2)$ theory, the dimension of the gauge group is ten, i.e., intermediate between those of $\operatorname{\mathrm{SU}}(3)$ and $\operatorname{\mathrm{SU}}(4)$, and a discontinuous deconfinement transition is observed; hence, if one would argue that this Yang–Mills theory should behave similarly to those based on unitary gauge groups, and would simply attempt to derive a crude estimate of the $L_{\mbox{\tiny{h}}}/T_{\mbox{\tiny{c}}}^{4}$ ratio by rescaling those found for the $\operatorname{\mathrm{SU}}(3)$ and $\operatorname{\mathrm{SU}}(4)$ theories according to the number of gluon degrees of freedom, one would get a result in the ballpark between approximately $1.5$ and $3$. The estimate (3.13) obtained from our simulation results lies within this interval.
@@
Figure 9: Extrapolation of our lattice results at $N_{\tau}=5$, $N_{\tau}=6$, and $N_{\tau}=7$ for the maximum value of the specific heat divided by the volume and expressed in the appropriate units of the lattice spacing (circles) to their infinite-volume limit. The dashed curves (for the $N_{\tau}=5$ and $N_{\tau}=6$ ensembles only) are the result of parabolic fits according to eq. (3.10), while the solid straight lines are obtained by linear fits of the data (i.e., setting $\gamma_{2}$ to zero). The corresponding extrapolated values, respectively denoted by the triangles and by the squares, lead to the values for the specific heat reported in the main text, which are compatible with each other.
@@
One of the reasons why it was interesting for us to study the finite-temperature confining transition in a symplectic Yang–Mills theory consisted in investigating, whether the center of the gauge group, which is widely believed to be directly responsible for the emergence of confinement at large distances in the ground state of non-Abelian gauge theories, via the formation of center vortices [157, 158, 159, 160], also affects the nature of the confinement/deconfinement transition at finite temperature. In particular, general arguments based on the renormalization group indicate that the thermal deconfinement transition for $\operatorname{\mathrm{Sp}}(N)$ gauge theories could be of second order, and in the same universality as the three-dimensional Ising model [67]. Our results, which are consistent with previous lattice studies of this issue [35], show that this is not the case: the transition is of first order, with a finite latent heat that, when expressed in units of the fourth power of the deconfinement temperature, is of the same order of magnitude as, and slightly larger than, in the $\operatorname{\mathrm{SU}}(3)$ Yang–Mills theory. This confirms the conjecture, that the continuous or discontinuous nature of the confinement/deconfinement transition is essentially determined by the imbalance in the number of degrees of freedom in the high- and in the low-temperature phases, i.e., it depends on the size of the gauge group, not directly on the center symmetry [35]. We also note that, in fact, the discontinuous nature of the deconfinement transition for the $\operatorname{\mathrm{Sp}}(2)$ theory is also supported by analytical arguments, including those based on calculations starting from a supersymmetric version of the theory on a manifold with a compactified spatial direction [221], and those derived within the framework of the functional renormalization group [222].
@@
4 Interpretation as a model for dark matter and gravitational-wave production
In this section we discuss some of the implications predicted from our model for the thermal evolution of dark matter in the early universe, and the features of the gravitational-wave power spectrum produced at the confining phase transition. Before doing so, however, an important remark is in order: if the dark sector is not directly coupled with the SM sector, then its temperature can be different from the temperature of the latter. In general, a difference in temperature among the visible and dark sectors could have been generated during reheating after inflation, if inflatons decay mostly into one of the two sectors, or if a sector contains a much larger number of heavy particle species than the other: when a thermalized particle species becomes non-relativistic, its entropy is transferred to the other particle species in the same sector (heating them up), eventually leading to a significant temperature imbalance between the two sectors [225].
If the dark sector is subject to reheating in the early stages of the universe, one expects that it will first be in the deconfined phase at high temperatures, then it will undergo a cosmological phase transition to the confining phase. Thus, in the model considered in this work, the non-Abelian gauge theory describing the dark sector will undergo an evolution qualitatively similar to the one of quantum chromodynamics, from a “hot”, deconfined phase, to a “cold”, confining phase.
Deconfined phase in the early universe. At sufficiently high temperatures realized in the early universe, the dark sector that we are studying consists of a plasma of $\operatorname{\mathrm{Sp}}(2)$ “dark gluons”, which interact with each other through screened long-range forces. Like in a generic non-Abelian gauge theory (see, for example, refs. [226, 227] and references therein), at sufficiently high temperatures a hierarchy of scales